OpenStudy (anonymous):

A plane travels from Orlando to Denver and back again. On the five-hour trip from Orlando to Denver, the plane has a tailwind of 40 miles per hour. On the return trip from Denver to Orlando, the plane faces a headwind of 40 miles per hour. This trip takes six hours. What is the speed of the airplane in still air?

OpenStudy (anonymous):

alright to you have any clue what to do first?

OpenStudy (anonymous):

Use d=r*t d is the same for both trip legs. 5 (r + 40) = 6 ( r - 40)